OpenStudy (anonymous):

3(x - 2) + 4(2x - 6) = 6(x - 4) + 8(2x + 1)

OpenStudy (anonymous):

You have to use distributive property.... LEts start with the left side. 3x - 6 + 8x - 24 then the right side. 8(2x + 1) 6x - 24 + 16x +8 so (3x -6) + (8x - 24) = (6x - 24) + (16x + 8) Then simplify. Add 3x and 8x from the left side. 11x. Then add -6 + (-24) = -30 so the left side should say 11x + (-30). The right side. 6x + 16x = 22x. Then add -24 + 8 = -16. So the right side shows 22x + (-16).

OpenStudy (anonymous):

so 11x + (-30) = 22x + (-16) Can you simplify this?

OpenStudy (anonymous):

i got -14/11

OpenStudy (anonymous):

Alrgiht what we have to do is add to both sides.... 11x - 30 = 22x - 16 +30 + 30 --------------------- 11x = 22x + 14 Then subtract 22 from 11. ---------------------- -11x = 14

OpenStudy (anonymous):

You have multiple choice?

OpenStudy (anonymous):

nope